OpenStudy (anonymous):

Please help! What is the slope of the line that crosses through the points (1,5) and (1,0)

jimthompson5910 (jim_thompson5910):

Two generic points that define a straight line are (x1, y1) and (x2, y2) In this case, (x1, y1) = (1, 5), which means x1=1 and y1=5 and (x2, y2) = (1, 0), which means x2=1 and y2=0 So x1=1, y1=5, x2=1, and y2=0. ------------------- Now use the slope formula m = (y2-y1)/(x2-x1) m = (0-5)/(1-1) m = -5/0 Since you cannot divide by zero, the slope is undefined.

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

A quick shortcut is this: if the x coordinates are the same, then you have a vertical line with an undefined slope.

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

cool thanks again

jimthompson5910 (jim_thompson5910):

sure thing